# Pattern formation of generalized fuzzy elementary cellular automaton

### v1

@@ Abstract
We propose a general method for constructing a fuzzy cellular automaton from a given cellular automaton. Unlike previous approaches that use fuzzy distinctive normal form, whose update function is restricted to third-order polynomials, our method accommodates a wide range of fuzzification functions, enabling the generation of diverse and complex time-evolution patterns that are unattainable with simpler heuristic models.
@@ 1 Introduction
A cellular automaton (CA) is a model of computation and a dynamical system that consists of a grid of cells, where each cell can take one of a finite number of states [1]. The system evolves in discrete time steps, with the state of each cell in the next step being determined by a fixed set of rules based on its current state and those of its neighboring cells. Even if its time evolution rule is simple, a CA can exhibit very complex behavior, generate a variety of patterns, and effectively capture how complex global patterns can arise from simple, local interactions. Hence, it is often used as a mathematical model for a natural or a social phenomenon [2].
An elementary cellular automaton (ECA) is the simplest possible type of CA [3]. It is a one-dimensional two-state three-neighboring CA, that is, the cells are arranged in a single line, each cell can only be in one of two states, typically represented as 0 (off/white) and 1 (on/black), and the state of a cell in the next generation is determined by its own state and the states of its immediate left and right neighbors from the current generation.
Despite its simplicity, it can generate very complex behavior, including fractal and chaotic patterns. It can be used as a mathematical model for pattern formation in chemical or biological systems. It also gives a fundamental mathematical model, a special case of the totally asymmetric simple exclusion process (TASEP), for traffic flow [4].
Although CAs have been used as models for natural and/or social phenomena, the state of a cell in a CA is limited to a discrete set, such as $\{0,1\}$, which has the advantage of being easy to implement on a computer. However, many real-world phenomena inherently involve uncertainty, ambiguity, or continuous quantities, which cannot be adequately represented by only two states. A fuzzy cellular automaton (FCA) is a system that attempts to overcome this limitation of the discrete model by extending the CA’s state space from a discrete set to a continuous interval through a process called fuzzification[5]. For a two-state CA (Boolean CA), a systematic method of fuzzification is to express the update rule in disjunctive normal form and replace it as [6]
@@
While it is almost impossible to predict the long-term behavior of a Boolean CA like an ECA, there have been reports of cases where mathematical analysis of global behavior and fixed points is possible for fuzzification of some ECAs [7, 8, 9, 10]. Furthermore, various applications of FCA to real-world systems have been actively discussed. Key areas include pattern recognition and information processing (e.g., noise removal, edge detection, and compression in image analysis; dataset generation via feature extraction in machine learning [11, 12, 13], and the modeling of complex systems (such as urban growth, traffic flow, and natural disasters like forest fires [14, 15, 16].
@@ 2 Fuzzification of ECAs via FDNF
Let us consider ECAs, which are one-dimensional Boolean CAs where the neighborhood of a cell consists of itself and its two immediate neighbors (a neighborhood size of three). There are $2^{8}=256$ possible time evolution rules for ECAs, conventionally numbered from 0 to 255. An ECA governed by rule $k$ is referred to as the rule $k$ ECA or ECA $k$ for short. When we denote by $u_{j}^{t}\in\{0,1\}$ ($j\in\mathbb{Z}$, $t\in\mathbb{Z}_{\geq 0}$) the value of the $j$ cite at time step $t$, the update rule of the rule $k$ ECA is expressed by a function $F_{k}:\,\{0,1\}^{\times 3}\rightarrow\{0,1\}$ as
@@
Now we consider the fuzzification of ECAs using FDNF [6]. From Eq. (1) and Table 1, the update function of an ECA is written in disjunctive normal form as
@@
Since $0\leq f_{k}(x,y,z)\leq 1$ for $(x,y,z)\in[0,1]^{3}$, and $f_{k}(x,y,z)=F_{k}(x,y,z)$ for $(x,y,z)\in\{0,1\}^{3}$, $f_{k}(x,y,z)$ serves as an update function of the fuzzy elementary ellular automaton (FECA). For instance, application of Eq. (3) to the rule 184 ECA yields
@@
Let us examine time evolution patterns of the FECAs. In contrast to the clear structure seen in Figure 1, the time evolution of rule 184 FECA (Figure 2) appears largely uniform. This is because the system is provably destined to become a monotone pattern when the number of cells is odd, and a checkerboard pattern when it is even[15].
@@ 3 Fuzzification function and generalized fuzzy elementary cellular automata
Let us remind that a Boolean FCA constructed from a CA should satisfy the two conditions: (i) the value of its cell is in $[0,1]$, and (ii) time evolution rule coincides with that of the CA when its initial values are in $\{0,1\}$. Hence we define
@@ 1. FECA converging to ECA behavior
Using Eq. (7), we obtain rule 90 and 184 GFECAs. Figures 7 and 7 show time evolution patterns of the rule 90 and 184 GFECAs respectively thus obtained. We used random initial values in the interval $[0,1]$ with periodic boundary conditions. Although initial values vary from 0 and 1, they eventually converge to binary values over time, reproducing the original ECA behavior. In the extreme case where $u(x)$ is defined as a step function:
@@ 2. Functions with a gap at $\boldsymbol{\frac{1}{2}}$
In the range $1.1\leq a\leq 1.6$, the patterns are notably similar to those of the standard rule 102 ECA (Fig. 9). For $a\geq 1.6$, this distinct structure disappears, giving way to a random-like pattern. We observe similar behavior in case of the rule 195 GFECA.
@@
Figure 10 shows the rule 184 GFECA obtained by the same fuzzification functions and parameters. For $a=1.0$, the system evolves into a uniform state as expected too. As $a$ increases, a labyrinthine pattern emerges around $1.1\leq a\leq 1.3$, and changes to a random-like pattern for $1.4\leq a$. The similar behavior is observed in the rule 226 GFECA.
@@ 5 FCAs by mixing of ECA rules
To quantify the complexity of the evolving patterns, we introduce an order parameter $d_{\alpha}$. This parameter is calculated for an analysis domain $D$ containing $M$ cells, where the state of a cell at position $(i,j)$ is given by $\phi_{ij}$.
@@ 6 GFECA dynamics with small degrees of freedom
Cellular automata with a small number of cells typically do not exhibit complex patterns. For example, an ECA with only three cells exhibits periodic time evolution over its $8$ possible states, with the period’s maximum length of $8$ depending on the boundary conditions. However, a GFECA can exhibit complex time evolution if its fuzzification functions are chosen properly. As an example, consider a GFECA with three cells where the update function Eq. (6) uses the identity function for $g(x)$, and the functions $u(x)$, $v(x)$, $w(x)$ are identical and given by
@@
The fuzzification function Eq. (13) is shown in Fig. 5 ($f$). Let $x_{t}$, $y_{t}$ and $z_{t}$ be the values of the three cells at time $t$. Then, the dynamics of rule $k$ GFECA are governed by the following system of equations:
@@
As shown in Fig. 15, the orbits of the GFECA in the cube $0,1]^{3}$ converges to a period-two orbit consisting of the points $(0,0,0)$ and $(1,1,1)$. The figure’s left panel plots the points from $t=0$ to $1000$, and the right panel plots the final stable orbit from $t=9000$ to $10000$. Similar behavior is observed in GFECA with rule $1$, 55, 77, and $17$.
@@
Other rules converge to longer cycles. For instance, rule $210$ converges to a period-six orbit (Fig. 16). Similar periodic convergence is observed in the GFECAs with rule $154$, $156$, and $198$.
@@
where $u(y)$ is the inherently chaotic fuzzification function from Eq. (13). Consequently, the GFECA under rule $85$ inherits this chaotic dynamics directly.
@@
This study focused on a three-cell GFECA—the minimal configuration for a three-neighbor FCA—using only the fuzzification function in Eq. (13). Increasing the number of cells would create dynamical systems with more degrees of freedom, likely leading to more complex evolutionary patterns and asymptotic behaviors. Furthermore, exploring a wider variety of fuzzification functions is a promising avenue for discovering richer and more diverse dynamical systems within the GFECA framework.
@@ 7 Concluding Remarks
This paper introduces a general method for constructing a FCA from any given Boolean CA. Since a multi-state CA can be replaced by equivalent Boolean CAs (for example, see [17]), the method is also applicable to multi-state CAs. Unlike the heuristic FDNF method, our approach not only retains the characteristic patterns of the original CA but also generates more complex patterns by using diverse fuzzification functions. We applied this method to ECAs, creating GFECAs. We demonstrated that by selecting the fuzzification function, these GFECAs exhibit controlled pattern transformations based on the function’s parameters. Furthermore, our method facilitates the mixing of ECA rules, which leads to observable phase transitions as the mixing ratio is altered. These transitions were quantitatively confirmed using the effective dimension related to $\ell^{p}$-norm. Notably, the resulting GFECAs display non-trivial temporal evolution even in minimal systems of just three cells. In essence, we have developed a general fuzzification method that preserves the complex behaviors of CAs, is mathematically tractable, and is expected to expand their utility as a modeling tool for natural and social phenomena.
Future research will focus on integrating our FCA framework with machine learning to explore new applications. This integration will enable the systematic and automatic discovery and calibration of fuzzy transition rules, for instance, by embedding FCAs within neural networks to learn transition functions or by informing deep learning architectures. By establishing an effective method for modeling localized interactions under uncertainty, this work could be applied to diverse fields such as materials science, biology, and sociology. Addressing these challenges and exploring these applications will be the primary focus of our future work.

### v2

@@ Abstract
We propose a general method for constructing a fuzzy cellular automaton from a given cellular automaton. Unlike previous approaches that use fuzzy disjunctive normal form, whose update function is restricted to third-order polynomials, our method accommodates a wide range of fuzzification functions, enabling the generation of diverse and complex time-evolution patterns that are unattainable with simpler heuristic models.
@@ 1 Introduction
A cellular automaton (CA) is a model of computation and a dynamical system that consists of a grid of cells, where each cell can take one of a finite number of states [vonNeumann1966]. The system evolves in discrete time steps, with the state of each cell in the next step being determined by a fixed set of rules based on its current state and those of its neighboring cells. Even if its time evolution rule is simple, a CA can exhibit very complex behavior, generate a variety of patterns, and effectively capture how complex global patterns can arise from simple, local interactions. Hence, it is often used as a mathematical model for a natural or a social phenomenon [Wolfram2002].
An elementary cellular automaton (ECA) is the simplest possible type of CA [Wolfram1986]. It is a one-dimensional two-state three-neighboring CA, that is, the cells are arranged in a single line, each cell can only be in one of two states, typically represented as 0 (off/white) and 1 (on/black), and the state of a cell in the next generation is determined by its own state and the states of its immediate left and right neighbors from the current generation.
Despite its simplicity, it can generate very complex behavior, including fractal and chaotic patterns. It can be used as a mathematical model for pattern formation in chemical or biological systems. It also gives a fundamental mathematical model, a special case of the totally asymmetric simple exclusion process (TASEP), for traffic flow [Nagel1992].
Although CAs have been used as models for natural and/or social phenomena, the state of a cell in a CA is limited to a discrete set, such as $\{0,1\}$, which has the advantage of being easy to implement on a computer. However, many real-world phenomena inherently involve uncertainty, ambiguity, or continuous quantities, which cannot be adequately represented by only two states. A fuzzy cellular automaton (FCA) is a system that attempts to overcome this limitation of the discrete model by extending the CA’s state space from a discrete set to a continuous interval through a process called fuzzification[Cattaneo1997]. For a two-state CA (Boolean CA), a systematic method of fuzzification is to express the update rule in disjunctive normal form and make the following replacements: [Betel2009]
@@
While it is almost impossible to predict the long-term behavior of a Boolean CA like an ECA, there have been reports of cases where mathematical analysis of global behavior and fixed points is possible for fuzzification of some ECAs [Flocchini2000, Mingarelli2003, Mingarelli2010, Yamamoto2023Asymptotic]. Furthermore, various applications of FCA to real-world systems have been actively discussed. Key areas include pattern recognition and information processing (e.g., noise removal, edge detection, and compression in image analysis; dataset generation via feature extraction in machine learning [Nayak2014, Sadjadi2022, Tangsakul2023], and the modeling of complex systems (such as urban growth, traffic flow, and natural disasters like forest fires [Foroutan2022, Higashi2021, Nishida2022].
@@ 2 Fuzzification of ECAs via FDNF
Let us consider ECAs, which are one-dimensional Boolean CAs where the neighborhood of a cell consists of itself and its two immediate neighbors (a neighborhood size of three). There are $2^{8}=256$ possible time evolution rules for ECAs, conventionally numbered from 0 to 255. An ECA governed by rule $k$ is referred to as the rule $k$ ECA or ECA $k$ for short. When we denote by $u_{j}^{t}\in\{0,1\}$ ($j\in\mathbb{Z}$, $t\in\mathbb{Z}_{\geq 0}$) the value of the $j$ cell at time step $t$, the update rule of the rule $k$ ECA is expressed by a function $F_{k}:\,\{0,1\}^{\times 3}\rightarrow\{0,1\}$ as
@@
Now we consider the fuzzification of ECAs using FDNF [Betel2009]. From Eq. (1) and Table 1, the update function of an ECA is written in disjunctive normal form as
@@
Since $0\leq f_{k}(x,y,z)\leq 1$ for $(x,y,z)\in[0,1]^{3}$, and $f_{k}(x,y,z)=F_{k}(x,y,z)$ for $(x,y,z)\in\{0,1\}^{3}$, $f_{k}(x,y,z)$ serves as an update function of the fuzzy elementary cellular automaton (FECA). For instance, application of Eq. (3) to the rule 184 ECA yields
@@
Let us examine time evolution patterns of the FECAs. In contrast to the clear structure seen in Figure 1, the time evolution of rule 184 FECA (Figure 2) appears largely uniform. This is because the system is provably destined to become a monotone pattern when the number of cells is odd, and a checkerboard pattern when it is even[Higashi2021].
@@ 3 Fuzzification function and generalized fuzzy elementary cellular automata
Let us recall that a Boolean FCA constructed from a CA should satisfy the two conditions: (i) the value of its cell is in $[0,1]$, and (ii) time evolution rule coincides with that of the CA when its initial values are in $\{0,1\}$. Hence we define
@@ 1. FECA converging to ECA behavior
Using Eq. (7), we obtain rule 90 and 184 GFECAs. Figures 7 and 7 show time evolution patterns of the rule 90 and 184 GFECAs respectively thus obtained. We used random initial values in the interval $[0,1]$ with periodic boundary conditions. Although the initial values are not binary (i.e., not just 0 and 1), they eventually converge to binary values over time, reproducing the original ECA behavior. In the extreme case where $u(x)$ is defined as a step function:
@@
To investigate how pattern formation is influenced by the derivative of the fuzzification function at the boundaries 0 and $1$, we introduce a one-parameter family of functions $u(x)$ defined as:
$u(x)=\begin{cases}ax&(0\leq x\leq\frac{1}{3-a})\\ x&(\frac{1}{3-a}<x\leq\frac{2-a}{3-a})\\ ax+(1-a)&(\frac{2-a}{3-a}<x\leq 1)\end{cases}$ (8)
Here, the parameter $a$ varies within the interval $[0,1]$. Specifically, $a=0$ recovers Eq. (7), while $a=1$ corresponds to the FDNF. The time evolution patterns for GFECA rules 30, 90, and 184, initiated from random values, are shown in Figs. 8, 9, and 10, respectively.
Figure 8: Time evolution patterns of the rule 30 GFECA using the fuzzification function in Eq. (8). The system size is 100 sites with periodic boundary conditions. The parameter $a$ takes values $0.0,0.2,0.4,0.6,0.8,\text{and }1.0$.
Figure 9: Time evolution patterns for the rule 90 GFECA. The parameters are identical to those in Fig. 8.
Figure 10: Time evolution patterns for the rule 184 GFECA. The parameters are identical to those in Fig. 8.
As the derivative $a$ at the boundaries approaches 1, the patterns become increasingly blurred; however, the degree of blurring is rule-dependent. To quantify this effect, we plot the difference between the maximum and minimum site values after 100 time steps in Figs. 11(a)–(c). These results imply that pattern dynamics depend significantly on the interplay between the specific fuzzification function and the underlying CA rules.
(a) Rule 30 GFECA
(b) Rule 90 GFECA
(c) Rule 184 GFECA
Figure 11: The horizontal axis represents the parameter $a$, against which we plot the difference between the maximum and minimum site values. The resulting behavior is highly rule-specific.
@@ 2. Functions with a gap at $\boldsymbol{\frac{1}{2}}$
In the range $1.1\leq a\leq 1.6$, the patterns are notably similar to those of the standard rule 102 ECA (Fig. 13). For $a\geq 1.6$, this distinct structure disappears, giving way to a random-like pattern. We observe similar behavior in the case of the rule 195 GFECA.
@@
Figure 14 shows the rule 184 GFECA obtained by the same fuzzification functions and parameters. For $a=1.0$, the system evolves into a uniform state as expected too. As $a$ increases, a labyrinthine pattern emerges around $1.1\leq a\leq 1.3$, and changes to a random-like pattern for $1.4\leq a$. Similar behavior is observed in the rule 226 GFECA.
@@ 5 FCAs by mixing of ECA rules
To quantify the complexity of the evolving patterns, we define an order parameter $d_{\alpha}$ over an analysis domain $D$ comprising $M$ cells. Here, $\phi_{ij}$ represents the state of the cell at position $(i,j)$. This order parameter is derived from the $\ell^{p}$ norm of a normalized function and is closely related to the multifractal dimension [Halsey1986].
@@ 6 GFECA dynamics with small degrees of freedom
Cellular automata with a small number of cells typically do not exhibit complex patterns. For example, an ECA with only three cells exhibits periodic time evolution over its $8$ possible states, with a maximum period length of 8, depending on the boundary conditions. However, a GFECA can exhibit complex time evolution if its fuzzification functions are chosen properly. As an example, consider a GFECA with three cells where the update function Eq. (5) uses the identity function for $g(x)$, and the functions $u(x)$, $v(x)$, $w(x)$ are identical and given by
@@
The profile of this fuzzification function is illustrated in Fig. 5(f). Notably, the one-dimensional mapping defined by Eq. (14), given by
$x_{n+1}=u(x_{n})\quad(0\leq x_{n}\leq 1),$ (15)
exhibits chaotic dynamics. Specifically, Eq. (15) possesses the following period-3 cycle:
$\frac{1}{14}\rightarrow\frac{3}{14}\rightarrow\frac{9}{14}\rightarrow\frac{1}{14}$
According to the Li-Yorke theorem ("Period Three Implies Chaos"), the existence of this cycle proves that the mapping in Eq. (15) is chaotic. Furthermore, since $|u^{\prime}(x)|=3$ for almost every $x\in[0,1]$, the Lyapunov exponent for a generic orbit is readily calculated as:
$\lambda=\lim_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n-1}\ln|u^{\prime}(x_{k})|=\ln 3>0$
Let $x_{t}$, $y_{t}$ and $z_{t}$ be the values of the three cells at time $t$. Then, the dynamics of rule $k$ GFECA are governed by the following system of equations:
@@
As shown in Fig. 19, the orbits of the GFECA in the cube $0,1]^{3}$ converges to a period-two orbit consisting of the points $(0,0,0)$ and $(1,1,1)$. The figure’s left panel plots the points from $t=0$ to $1000$, and the right panel plots the final stable orbit from $t=9000$ to $10000$. Similar behavior is observed in GFECA with rules $1$, 55, 77, and $17$.
@@
Other rules converge to longer cycles. For instance, rule $210$ converges to a period-six orbit (Fig. 20). Similar periodic convergence is observed in the GFECAs with rules $154$, $156$, and $198$.
@@
We now investigate the orbital characteristics of rule 210 GFECA. The transition function $\tilde{f}_{210}(x,y,z)$ possesses the following symmetry properties:
${}^{\forall}y,\,{}^{\forall}z,\ \tilde{f}_{210}(1/2,y,z)$ $=1/2$ (17)
${}^{\forall}x,\,{}^{\forall}y,\,{}^{\forall}z,\ \tilde{f}_{210}(x,y,z)$ $=1-\tilde{f}_{210}(1-x,1-z,1-y)$ (18)
From (17), we identify a specific orbit of the form:
$(x_{0},1/2,1/2)\rightarrow(1/2,x_{1},1/2)\rightarrow(1/2,1/2,x_{2})\rightarrow x_{3},1/2,1/2)\rightarrow(1/2,x_{4},1/2)\rightarrow\cdots$
where the evolution is governed by the one-dimensional map:
$x_{n+1}=\frac{1}{4}+\frac{1}{2}u(x_{n})$
As the Lyapunov exponent for a generic orbit of this map is $\log\frac{3}{2}>0$, this specific subspace exhibits chaotic behavior.
However, numerical simulations indicate that a significant portion of initial states in the full 3D space converges to one of two period-6 cycles. By utilizing the symmetry properties described above, algebraic analysis reveals the first period-6 cycle:
$\cdots\rightarrow(a,1/2,b)\rightarrow(c,d,1/2)\rightarrow(1/2,b,a)\rightarrow(d,1/2,c)\rightarrow(b,a,1/2)\rightarrow(1/2,c,d)\rightarrow\cdots$
and its symmetric counterpart:
$\cdots\rightarrow(\bar{b},1/2,\bar{a})\rightarrow(\bar{d},\bar{c},1/2)\rightarrow(1/2,\bar{a},\bar{b})\rightarrow(\bar{c},1/2,\bar{d})\rightarrow(\bar{a},\bar{b},1/2)\rightarrow(1/2,\bar{d},\bar{c})\rightarrow\cdots$
where $\bar{x}:=1-x$. The values $a\simeq 0.478535$, $b\simeq 0.623557$, $c\simeq 0.290795$, and $d\simeq 0.556067$ correspond to specific roots of the following cubic equations, respectively:
$t^{3}-\frac{67}{2\cdot 3\cdot 5}t^{2}+\frac{7^{2}\cdot 113}{2\cdot 3^{3}\cdot 5\cdot 13}t-\frac{16729}{3^{6}\cdot 5\cdot 13}$ $=0$
$t^{3}-\frac{7}{2\cdot 3}t^{2}+\frac{107}{2\cdot 3^{3}\cdot 5}t-\frac{131}{3^{6}\cdot 5}$ $=0$
$t^{3}-\frac{5}{2\cdot 3}t^{2}+\frac{151}{2\cdot 3^{3}\cdot 13}t-\frac{2\cdot 79}{3^{6}\cdot 13}$ $=0$
$t^{3}-\frac{47}{2\cdot 3\cdot 5}t^{2}+\frac{191}{2\cdot 3^{3}\cdot 5}t-\frac{2\cdot 11\cdot 67}{3^{6}\cdot 5^{2}}$ $=0$
Each equation yields three real roots; $a$ is the smallest root, $b$ is the largest, while $c$ and $d$ are the intermediate roots of their respective equations.
[…]
Figure 21: Basins of attraction for the two period-6 cycles of Rule 210 GFECA on the slice $z=1/2$. The left and middle panels show the individual basins for the first and second cycles, respectively. The right panel shows the superposition of both basins. Black denotes convergence to the first cycle, and gray denotes convergence to the second.
Figure 22: Basins of attraction for the slice at $z=0.3$. Compared to the $z=1/2$ slice, the basins here exhibit a highly intermingled, dust-like structure.
@@
where 1u(y)1-u(y) is the inherently chaotic fuzzification function just like the function Eq. (14), because it has a period-3 cycle:
$\frac{1}{28}\rightarrow\frac{25}{28}\rightarrow\frac{9}{28}\rightarrow\frac{1}{28}$
Consequently, the GFECA under rule $85$ inherits this chaotic dynamics directly.
@@
This study focused on a three-cell GFECA—the minimal configuration for a three-neighbor FCA—using only the fuzzification function in Eq. (14). Although the fuzzification function itself gives a typical chaotic mapping, the three-cell GFECAs show various dynamics depending on the ECA rules. Increasing the number of cells would create dynamical systems with more degrees of freedom, likely leading to more complex evolutionary patterns and asymptotic behaviors. Furthermore, exploring a wider variety of fuzzification functions is a promising avenue for discovering richer and more diverse dynamical systems within the GFECA framework.
@@ 7 Concluding Remarks
This paper presents a generalized method for constructing Fuzzy Cellular Automata (FCA) from any given Boolean CA. Given that multi-state CAs can be represented by equivalent Boolean frameworks (e.g., [Nishida2020]), our method is also applicable to multi-state systems. Unlike the heuristic Fuzzy Disjunctive Normal Form (FDNF) approach, our method not only preserves the characteristic patterns of the original CA but also generates more complex behaviors through various fuzzification functions.
We applied this framework to ECAs to develop Generalized Fuzzy Elementary Cellular Automata (GFECAs). Our results demonstrate that selecting specific fuzzification functions enables controlled pattern transformations governed by the functions’ parameters. Furthermore, the proposed method facilitates the mixing of ECA rules, leading to observable phase transitions as the mixing ratio is varied. These transitions were quantitatively verified using the effective dimension derived from the $\ell^{p}$-norm. Notably, the resulting GFECAs exhibit non-trivial temporal evolution even in minimal systems of only three cells.
A primary objective for future research is to systematize these findings into a framework of general properties. Since the dynamics are highly sensitive to both the GFECA structure and the specific functions employed, a comprehensive classification requires further extensive investigation, which will be addressed in future works.
In essence, we have developed a mathematically tractable fuzzification method that preserves the complex behaviors of CAs, thereby expanding their utility as modeling tools for natural and social phenomena. Future research will explore the integration of our FCA framework with machine learning. This synergy may enable the systematic discovery and calibration of fuzzy transition rules—for instance, by embedding FCAs within neural networks to learn transition functions or by using them to inform deep learning architectures. By providing an effective method for modeling localized interactions under uncertainty, this work holds potential for diverse fields, including materials science, biology, and sociology.